OpenStudy (anonymous):

Solve the quadratic equation. (x + 2)2 = 1 A) x = 1 or 3 B) x = -1 or -3 C) x = ± 3 D) x = ± 1

OpenStudy (saifoo.khan):

r u sure this is the correct ques?

OpenStudy (anonymous):

yes

OpenStudy (gw2011):

The answer is B where x = -1 or -3 Here is how you get the answer: (x+2)^2 = 1 Take the square root of both sides and you get: x+2 = +/- square root (1) = +/- 1 x = -2 + 1 = -1 x = -2 - 1 = -3